OpenStudy (anonymous):

Evaluate 3x^2-2xy+4y^2 for x=-3 y=2

OpenStudy (callisto):

Put x=-3 and y=2 into 3x^2-2xy+4y^2 3(-3)^2-2(-3)(2)+4(2)^2 = ...?

OpenStudy (anonymous):

i have27+6-4+16=-21

OpenStudy (anonymous):

81+12+64= 157

OpenStudy (anonymous):

that answer is not right

OpenStudy (callisto):

3(-3)^2-2(-3)(2)+4(2)^2 = 3(9) - (-6)(2) + 4(4) = 27 + 12 + 16 = ...?

OpenStudy (anonymous):

55

OpenStudy (anonymous):

i got 31??

OpenStudy (callisto):

@katiebugg Can you show your workings??

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

the answer is 31

OpenStudy (anonymous):

27-12+16=31

OpenStudy (anonymous):

told you lol

OpenStudy (callisto):

No. You made a mistake or the question is wrong

OpenStudy (callisto):

For the middle term -2(-3)(2) = -2 ( -6) = 12 negative times negative gives positive.

OpenStudy (anonymous):

3(9)-2(-3)(2)+4(4) 27+12+16=55 i did i allways mess up the signs when i do it in my head :( sorry i need to stop that

OpenStudy (callisto):

@kaleeb20 55 is right

OpenStudy (anonymous):

ok thank you guys

OpenStudy (anonymous):

yeah idk y ur thankin us lol u did it!!

OpenStudy (anonymous):

lol just trying to be nice dats all lol

OpenStudy (anonymous):

i know lol thanks!!